I have the sequence 2,5,9,14,20,27,35,44...
How can I find an explicit formula to find the nth term? How can I find the nth term if An = 324?

Found 2 solutions by stanbon, 0123456:
Answer by stanbon(75887)

I have the sequence 2,5,9,14,20,27,35,44...
How can I find an explicit formula to find the nth term? How can I find the nth term if An = 324?
---------------
a(1) = 2
a(2) = a(1)+3
a(3) = a(2)+4
a(4) = a(3)+5
....
a(n) = a(n-1)+(n+1)
------------------------
Cheers,
Stan H.

Answer by 0123456(1)

Общий член:
An=(n^2+3*n)/2